# FP4DiT: Towards Effective Floating Point Quantization for Diffusion Transformers

### v1

@@ 3.3 Token-wise Activation Quantization
Prior DM PTQ approaches [11, 15] use a calibration dataset to learn temporally-aware [10] activation quantization scales. This approach is predicated on knowledge of how U-Net activation distributions change as a function of denoising timesteps, i.e., activation ranges taper-off towards the end of the denoising process [45, 23].
In contrast, we show that this assumption does not hold for DiT models. First, we collect input activations of PixArt-$\alpha$ across 20 timesteps, revealing that the activation range remains stable over time, as shown in Figure 5(a). We then analyze the Adaptive Layer Normalization (AdaLN) in the PixArt DiT model in Figure 2. Since the feed-forward scale directly influences the DiT block output range, we visualize the feed-forward scale in the first layer in Figure 5(b) and the output of the 7th DiT block in Figure 5(c). These figures demonstrate that the value of activations is primarily controlled by channels opposed to timesteps, and that the width of the activation range tends to remain constant, but shifts as a function of time.
@@
Table 1: CLIP score [12] reported when quantizing PixArt-$\alpha$ to 4-bit weights (W4) and 8 or 6-bit activations (A8 and A6) using the online token-wise method and temporarally-aware scale calibration. Specifically, we generate 1k images per configuration using COCO [26] prompts and compare against the validation set. Higher CLIP means better prompt-image alignment.
Recent works by Microsoft [57, 53] suggest that token-wise online activation quantization approaches may yield superior results when quantizing transformer activations, especially when paired with kernel fusion [51] techniques. Table 1 applies this hypothesis to the DiT scenario. Specifically, we apply simple min-max quantization to reduce weight precision of PixArt-$\alpha$ to 4-bits (W4), then consider 8-bit (A8) and 6-bit (A6) activation quantization. In both scenarios, we observe CLIP performance that is closer to the full precision model using token-wise activation quantization as opposed to the traditional, temporally-aware scale calibration technique which is designed for U-Nets. As such, we consider token-wise activation quantization throughout the remainder of this work by substituting it into U-Net baselines like Q-Diffusion [23] and TFMQ-DM [15].
4 Results
In this section we conduct experiments to verify the efficacy of FP4DiT. We elaborate on our experimental setup and then compare FP4DiT to several baselines approaches to highlight its competitiveness in terms of quantitative metrics and qualitative image generation output. Specifically, we consider three text-to-image (T2I) models: PixArt-$\alpha$ [3], PixArt-$\Sigma$ [2] and Hunyuan [25]. We also conduct several ablation studies to verify the components of our method. Finally, we report several hardware metrics tabulating the cost savings and throughput of FP4DiT.
@@ 4.1 Experimental Settings
We use the HuggingFace Diffusers library [50] to instantiate the base DiT models in W16A16 bit-precision and consider the default values for inference parameters like number of denoising steps and classifier-free guidance (CFG) scale. We quantize weights to 4-bit precision FP format. Specifically, we set the weight format for the first linear layer in each pointwise feed-forward to be E3M0. We quantize all other weights to E2M1 for PixArt-$\alpha$ and Hunyuan, and E1M2 for PixArt-$\Sigma$. Further details on this decision are provided in the Supplementary Materials. Finally, our weight quantization is group-wise [9, 40] along the output channel dimension with a group size of 128.
We perform weight quantization calibration using our scale-aware AdaRound and BRECQ. Weight calibration requires a small amount of calibration data. We use 128 (64 for Hunyuan) image-text pairs from the MS-COCO 2017 train [26] dataset and calibrate for 2.5k iterations per DiT block or layer. Next, we perform activation quantization to 8 or 6-bit precision using min-max token-wise quantization from ZeroQuant [57, 53]. We provide further hyperparameter details in the Supplementary Materials.
Model Method Precision HPSv2 $\uparrow$ CLIP $\uparrow$
PixArt-$\alpha$ Full Precision W16A16 31.01 0.3075
Q-Diffusion W4A8 23.17 0.3017
TFMQ-DM W4A8 25.99 0.3066
ViDiT-Q W4A8 17.42 0.2900
FP4DiT (ours) W4A8 27.43 0.3076
Q-Diffusion W4A6 12.5 0.2868
TFMQ-DM W4A6 23.29 0.3015
ViDiT-Q W4A6 16.44 0.2827
FP4DiT (ours) W4A6 23.55 0.3031
PixArt-$\Sigma$ Full Precision W16A16 31.67 0.3139
Q-Diffusion W4A8 26.13 0.3050
TFMQ-DM W4A8 23.62 0.2975
ViDiT-Q W4A8 26.07 0.2562
FP4DiT (ours) W4A8 26.27 0.3064
Q-Diffusion W4A6 22.67 0.3027
TFMQ-DM W4A6 18.53 0.2346
ViDiT-Q W4A6 23.37 0.2425
FP4DiT (ours) W4A6 25.40 0.3040
Hunyuan Full Precision W16A16 32.08 0.3102
Q-Diffusion W4A8 24.96 0.3006
TFMQ-DM W4A8 27.56 0.3075
FP4DiT (ours) W4A8 27.78 0.3102
Q-Diffusion W4A6 14.83 0.2277
TFMQ-DM W4A6 13.68 0.2520
FP4DiT (ours) W4A6 15.00 0.2562
Table 2: Quantitative evaluation results for PixArt-$\alpha$, PixArt-$\Sigma$ and Hunyuan in terms of average score on the HPSv2 benchmark and average CLIP score using 10k (3k for Hunyuan) COCO 2017 prompt-image pairs. Best and second best results in bold and italics, respectively.
@@ 4.2 Main Results
Figure 7: PixArt-$\alpha$ (up) and PixArt-$\Sigma$ (down) images and comparison between FP4DiT and related work. Best viewed in color.
Figure 8: Hunyuan image comparison. Note the ‘white background’ detail for FP4DiT. Best viewed in color.
In our experiment, we consider three baseline approaches: Q-Diffusion [23], TFMQ-DM [15] and ViDiT-Q [58]. Note that while Q-Diffusion and TFMQ-DM are originally designed for U-Nets, we modify these approaches to use the same online token-wise activation quantization as FP4DiT per Table 1, while ViDiT-Q uses this mechanism by default. Further, note that ViDiT-Q uses mix-precision meaning some of their layers are not quantized to 4-bits. For the sake of convenience, we use their mix-precision as a W4 baseline to conduct our experiments.
We generate $512\times 512$ resolution images using PixArt-$\alpha$ and $1024\times 1024$ for PixArt-$\Sigma$ and Hunyuan. For evaluation, we primarily consider the Human Preference Score v2 (HPSv2) [52] benchmark. Specifically, HPSv2 considers four image categories: animation, concept-art, painting and photography, and estimates the human preference score of an image generated using a prompt with respect to one category. Each category contains 800 prompts, requiring 3.2k images be generated to fully evaluate. The final HPSv2 score is the average estimated human preference across all four categories. Additionally, we measure the CLIP [12] score using prompts from the MS-COCO 2017 [26] validation set.
Table 2 compares our method with the stated baseline approaches at the W4A8 and W4A6 precision levels. FP4DiT consistently outperforms all other methods across three different base models at every precision level. Specifically, for the PixArt models, W4A6 precision causes the baselines to experience a significant performance drop, while FP4DiT still maintains a balanced trade-off between accuracy and efficiency. Also, results on Hunyuan further demonstrate the efficacy of our method at A8 precision, and although A6 quantization is substantially difficult, FP4DiT still provides the best results. Further, due to space constraints, we report performance on each individual HPSv2 category as well as other metrics [13] in the in the supplementary. Overall though, Table 2 demonstrates the efficacy of FP4DiT in term of quantitative T2I compared to prominant baseline approaches.
@@
Further, Figure 8 provides image results on Hunyuan, where we again note the detail present in the FP4DiT image, while the baseline approaches are much noisier and have yellow backgrounds which are not prompt-adherent. Finally, additional visualization results can be found in the Supplementary Materials.
@@ 4.2.1 User Preference Study
We conduct a preference study to qualitatively compare FP4DiT to baseline approaches. Specifically, we consider 75 random prompts and 15 human participants. Each participant is presented with a prompt and the corresponding image generated by Hunyuan when quantized to W4A8 or W4A6 precision by FP4DiT, Q-Diffusion and TFMQ-DM. We record the percentage that each approach is selected as the preferred image and report the results in Table 3. At A8 precision FP4DiT clearly outperforms the other methods with a majority of the images being favored, while it also obtains over one third of the votes at A6 precision.
4.3 Ablation Studies
We conduct ablation studies on the PixArt-$\alpha$ model to verify the contribution of each component of FP4DiT. The experiment settings are consistent with Section 4.2 unless specified. Additional ablation studies can be found in the Supplementary Materials.
Model Method Precision Preference(%) $\uparrow$
@@ 4.3 Ablation Studies
FP4DiT (ours) W4A6 38.81
Table 3: User preference study between FP4DiT and baseline methods on Hunyuan DiT with W4A8 and W4A6 quantization.
4.3.1 Effect of Weight Quantization
@@ 4.3.2 Effect of Scale-Aware AdaRound
To further verify our scale-aware AdaRound for FP quantization, we compare our scale-aware AdaRound to the origin AdaRound with INT quantization in the W4A16 setting. Note that the calibration budget is crucial for the performance of weight reconstruction. BRECQ [24] uses 20k as default and this setting is inherited by prior DM quantization research like Q-Diffusion and TFMQ-DM. Thus, we configured calibration budgets at {2.5k, 5k, 10k, 20k} to ensure a fair comparison. Figure 9 outlines the CLIP of different AdaRound methods on PixArt-$\alpha$ under different calibration budgets. Scale-aware AdaRound achieving optimal budget with 8 times fewer calibration steps than INT AdaRound, highlights its effectiveness in reducing calibration costs without compromising reconstruction quality.
Figure 9: The performance of different AdaRound methods with different calibration budgets on W4A16 PixArt-$\alpha$ quantization generating 1k $512\times 512$ images using COCO prompts. The stars indicate the optimal budget for INT and scale-aware AdaRound. The scale-aware AdaRound achieves an 8-fold reduction in calibration computation cost.
@@ 4.4 Hardware Cost Comparison
We compare the hardware cost of the quantized FP4DiT model with the full-precision model using two metrics: model size and Bit-Ops (BOPs) [11]. Specifically, model size is the disk space required to store the model checkpoint weights and scales, while BOPs is a quantization-aware extension of the MACs [4, 32] metric which measures the compute cost of a neural network forward pass.
Table 5 outlines our results. We see that the group-wise weight quantization only brings moderate overhead hence our method still substantially reduces the model size and BOPs computational complexity.
Finally, in terms of computational throughput, Table 6 shows the computation rate of different INT and FP formats across different Nvidia GPUs. Although the floating-point math is typically more complex, the state-of-the-art Nvidia “Blackwell” GPUs like the 5090, B100 and B200 are designed to handle INT and FP at the same computational throughput as shown, which ensures that FP-based quantization does not introduce additional computational overhead. Leveraging this, our FPQ method can achieve superior performance without sacrificing computational efficiency compared to INT quantization, making it an effective alternative. Although FP6 shares the same compute throughput as FP8, it brings memory saving (25% smaller tensor) which is essential for low-bit quantization, as transformers often become memory-bound [54] during token generation. As a result, FP6 enhances DiT’s inference efficiency.
@@
Table 5: The comparison of model size and latency of different Precision on PixArt-$\alpha$. G128 denotes group-wise weight quantization with a group size of 128. Lower is better.
GPU INT8 FP8 FP6 FP4
RTX 4090 [37] 660.6 TOPS 660.6 TFLOPS
H100 [39] 1979 TOPS 1979 TFLOPS
RTX 5090 [37] 838 TOPS 838 TFLOPS 838 TFLOPS 1676 TFLOPS
HGX B100 [38] 56 POPS 56 PFLOPS 56 PFLOPS 112 PFLOPS
HGX B200 [38] 72 POPS 72 PFLOPS 72 PFLOPS 144 PFLOPS
Table 6: GPU throughput rates for different low-bit datatype formats. Horizontal line demarcates older Ada Lovelace/Hopper GPUs from state-of-the-art Blackwell series. Older series don’t support FP6 and FP4.

### v2

@@ 3.3 Token-wise Activation Quantization
Prior DM PTQ approaches (He et al., 2024; Huang et al., 2024) use a calibration dataset to learn temporally-aware (He et al., 2023) activation quantization scales. This approach is predicated on knowledge of how U-Net activation distributions change as a function of denoising timesteps, i.e., activation ranges taper-off towards the end of the denoising process (Shang et al., 2023b; Li et al., 2023).
In contrast, we show that this assumption does not hold for DiT models. First, we collect input activations of PixArt-$\alpha$ across 20 timesteps, revealing that the activation range remains stable over time, as shown in Figure 5(a). We then analyze the Adaptive Layernorm (AdaLN) (Perez et al., 2018) in the PixArt DiT model in Figure 3. Since the feed-forward scale directly influences the DiT block output range, we visualize the feed-forward scale in the first layer in Figure 5(b) and the output of the 7th DiT block in Figure 5(c). These figures demonstrate that the value of activations is primarily controlled by channels opposed to timesteps, and that the width of the activation range tends to remain constant, but shifts as a function of time.
Recent works by Microsoft (Yao et al., 2022; Wu et al., 2023b) propose an online token-wise activation quantization that yields superior results when quantizing transformer activations. Table 1 applies this technique to the DiT scenario. Specifically, we apply simple min-max quantization to reduce weight precision of PixArt-$\alpha$ to 4-bits (W4), then consider 8-bit (A8) and 6-bit (A6) activation quantization. In both scenarios, we observe CLIP performance that is closer to the full precision model using token-wise activation quantization as opposed to the traditional, temporally-aware scale calibration technique which is designed for U-Nets. As such, we consider token-wise activation quantization throughout the remainder of this work by substituting it into U-Net baselines like Q-Diffusion (Li et al., 2023) and TFMQ-DM (Huang et al., 2024).
4 Results
@@
Table 1: CLIP score (Hessel et al., 2021) reported when quantizing PixArt-$\alpha$ to 4-bit weights (W4) and 8 or 6-bit activations (A8 and A6) using the online token-wise method and temporarally-aware scale calibration. Specifically, we generate 1k images per configuration using COCO (Lin et al., 2014a) prompts and compare against the validation set. Higher CLIP is better.
In this section we conduct experiments to verify the efficacy of FP4DiT. We elaborate on our experimental setup and then compare FP4DiT to several baselines approaches to highlight its competitiveness in terms of quantitative metrics and qualitative image generation output. Specifically, we consider three text-to-image (T2I) models: PixArt-$\alpha$ (Chen et al., 2024b), PixArt-$\Sigma$ (Chen et al., 2024a) and Hunyuan (Li et al., 2024). We also conduct several ablation studies to verify the components of our method. Finally, we report several hardware metrics tabulating the cost savings and throughput of FP4DiT.
@@ 4.1 Experimental Settings
We use the HuggingFace Diffusers library (von Platen et al., 2022) to instantiate the base DiT models in W16A16 bit-precision and consider the default values for inference parameters like number of denoising steps and classifier-free guidance (CFG) scale. We quantize weights to 4-bit precision FP format. Specifically, we set the weight format for the first linear layer in each pointwise feed-forward to be E3M0. We quantize all other weights to E2M1 for PixArt-$\alpha$ and Hunyuan, and E1M2 for PixArt-$\Sigma$. Further details on this decision are provided in Session 4.3.1. Finally, our weight quantization is group-wise (Frantar et al., 2022; Park et al., 2022) along the output channel dimension with a group size of 128.
We perform weight quantization calibration using our scale-aware AdaRound and BRECQ. Weight calibration requires a small amount of calibration data. We use 128 (64 for Hunyuan) prompts from the MS-COCO 2014 train (Lin et al., 2014a) dataset and calibrate for 2.5k iterations per DiT block or layer. Next, we perform activation quantization to 8 or 6-bit precision using min-max token-wise quantization from ZeroQuant (Yao et al., 2022; Wu et al., 2023b). We provide further hyperparameter details in the appendix.
Benchmark HPSv2 MS-COCO
[…]
PixArt-$\alpha$ Full Precision W16A16 32.56 31.06 30.76 29.67 31.01 34.05 0.3075
Q-Diffusion W4A8 24.18 23.43 22.91 22.15 23.17 52.20 0.3017
TFMQ-DM W4A8 27.88 26.03 25.14 24.91 25.99 64.73 0.3066
ViDiT-Q W4A8 17.93 17.35 16.81 17.59 17.42 39.11 0.2900
Q-DiT W4A8 28.49 26.91 27.55 26.56 27.38 36.17 0.3062
FP4DiT (ours) W4A8 28.63 26.79 27.59 26.72 27.43 30.37 0.3076
Q-Diffusion W4A6 12.63 13.39 13.32 10.65 12.50 70.96 0.2868
TFMQ-DM W4A6 24.02 23.36 22.72 23.04 23.29 90.09 0.3015
ViDiT-Q W4A6 16.71 16.28 16.23 16.56 16.44 56.54 0.2827
Q-DiT W4A6 24.66 22.59 23.29 23.04 23.40 43.91 0.3015
FP4DiT (ours) W4A6 24.57 23.08 23.30 23.26 23.55 40.61 0.3031
PixArt-$\Sigma$ Full Precision W16A16 33.07 31.58 31.54 30.49 31.67 36.94 0.3139
Q-Diffusion W4A8 27.30 26.11 26.06 25.06 26.13 36.89 0.3050
TFMQ-DM W4A8 24.95 23.59 23.19 22.73 23.62 62.98 0.2975
ViDiT-Q W4A8 27.10 26.45 26.24 24.49 26.07 37.17 0.2562
Q-DiT W4A8 27.74 26.23 26.05 24.67 26.17 32.03 0.3048
FP4DiT (ours) W4A8 27.95 26.29 26.16 24.67 26.27 31.58 0.3064
Q-Diffusion W4A6 24.07 22.42 22.47 21.72 22.67 74.83 0.3027
TFMQ-DM W4A6 19.30 18.46 18.85 17.50 18.53 154.13 0.2346
ViDiT-Q W4A6 24.84 23.23 23.48 21.94 23.37 87.47 0.2425
Q-DiT W4A6 25.67 23.57 23.60 22.64 23.87 73.20 0.2894
FP4DiT (ours) W4A6 26.91 25.55 25.22 23.93 25.40 42.21 0.3040
Hunyuan Full Precision W16A16 33.72 31.84 31.52 31.24 32.08 59.08 0.3102
Q-Diffusion W4A8 26.00 24.74 24.84 24.26 24.96 82.43 0.3006
TFMQ-DM W4A8 28.77 27.63 27.67 26.15 27.56 89.39 0.3075
ViDiT-Q W4A8 28.74 26.79 26.49 27.20 27.31 82.22 0.3099
Q-DiT W4A8 28.91 27.27 27.15 26.72 27.52 85.33 0.3088
FP4DiT (ours) W4A8 28.96 27.75 27.47 26.94 27.78 81.94 0.3102
Q-Diffusion W4A6 14.90 14.83 15.32 14.24 14.83 255.69 0.2277
TFMQ-DM W4A6 13.73 13.31 13.16 14.51 13.68 258.02 0.2520
ViDiT-Q W4A6 15.11 15.20 15.41 14.17 14.97 265.52 0.2559
Q-DiT W4A6 13.41 13.15 13.48 12.94 13.24 262.31 0.2312
FP4DiT (ours) W4A6 15.23 14.94 15.57 14.27 15.00 255.06 0.2562
Table 2: Quantitative evaluation results for PixArt-$\alpha$, PixArt-$\Sigma$ and Hunyuan in terms of HPSv2, FID, and CLIP score. Specifically, for each configuration, we generate 10k images (5k for Hunyuan) using COCO 2014 validation set prompts. Best and second best results in bold and italics, respectively.
@@ 4.2 Main Results
In our experiment, we consider four baseline approaches: Q-Diffusion (Li et al., 2023), TFMQ-DM (Huang et al., 2024), ViDiT-Q (Zhao et al., 2024) and Q-DiT (Chen et al., 2025). Note that while Q-Diffusion and TFMQ-DM are originally designed for U-Nets, we modify these approaches to use the same online token-wise activation quantization as FP4DiT per Table 1, while ViDiT-Q and Q-DiT uses this mechanism by default. Further, note that ViDiT-Q uses mix-precision meaning some of their layers are not quantized to 4-bits. For the sake of convenience, we use their mix-precision as a W4 baseline to conduct our experiments.
We generate $512\times 512$ resolution images using PixArt-$\alpha$ and $1024\times 1024$ for PixArt-$\Sigma$ and Hunyuan. For evaluation, we primarily consider the Human Preference Score v2 (HPSv2) (Wu et al., 2023a) benchmark and MS-COCO 2014 (Lin et al., 2014a) validation set. Specifically, HPSv2 considers four image categories: animation, concept-art, painting and photography, and estimates the human preference score of an image generated using a prompt with respect to one category. Each category contains 800 prompts, requiring 3.2k images to be generated to fully evaluate. The final HPSv2 score is the average estimated human preference across all four categories. Additionally, for COCO 2014, we measure the FID (Heusel et al., 2017) and CLIP (Hessel et al., 2021) score using prompts from the validation set.
Table 2 compares our method with the stated baseline approaches at the W4A8 and W4A6 precision levels. FP4DiT consistently outperforms all other methods across three different base models at every precision level in terms of each performance metric. On HPSv2, FP4DiT achieves the best average performance across every model and bit-precision combination. Although other approaches may achieve slightly higher performance on an individual category, these gaps are small, while FP4DiT leads in terms of FID and CLIP on the COCO 2014 validation set.
Next, we compare performance across activation bit-precision. For the PixArt-series DiT models, A6 precision can cause substantial performance degradation for some baselines, while FP4DiT still maintains high performance. This is specifically noteworthy on PixArt-$\Sigma$ as baseline approaches suffer a loss in terms of either FID and/or CLIP performance at A6 compared to A8, but this is not the case for FP4DiT. For the Hunyuan DiT model, A6 quantization remains highly challenging across the board, however, our method still achives the best results at this level. Overall though, Table 2 demonstrates the efficacy of FP4DiT in term of quantitative T2I compared to prominant baseline approaches.
@@
Further, Figure 8 provides image results on Hunyuan, where we again note the detail present in the FP4DiT image, while the baseline approaches are much noisier and have yellow backgrounds which are not prompt-adherent. Finally, additional visualization results can be found in the appendix.
Figure 7: PixArt-$\alpha$ (up) and PixArt-$\Sigma$ (down) images and comparison between FP4DiT and related work. Best viewed in color and zoomed in.
Figure 8: Hunyuan image comparison. Note the details like ‘white background’ and ‘detailed hair texture’ for FP4DiT. Best viewed in color.
@@ 4.2.1 User Preference Study
To further verify the utility of our method, we conduct several human user preference studies to qualitatively compare FP4DiT to existing baseline approaches. Specifically, for each human user preference study, we have a set of prompts (i.e., introduced in (Chen et al., 2024b)) as well as quantized variants of a single model (e.g., Hunyuan-DiT quantized to W4A8 using ViDiT-Q, Q-DiT and FP4DiT). Each quantized model generates one image per prompt. For each prompt, we solicit the opinion of a human participant, by presenting them the set of images produced using the prompt by different methods, and ask them to select which image they prefer, in terms of perceived prompt adherence and general visual quality. We then tally up the number of votes each method obtains and compute how often it is selected compared to its competitors, e.g., preference score (%), where higher is better.
Model Method Prec. Preference(%) $\uparrow$
PixArt-$\Sigma$ ViDiT-Q W4A6 17.09
Q-DiT W4A6 30.77
FP4DiT W4A6 52.14
Hunyuan ViDiT-Q W4A8 19.17
Q-DiT W4A8 35.83
FP4DiT W4A8 45.00
Table 3: DiT PTQ user preference study between FP4DiT, ViDiT-Q, and Q-DiT on PixArt-$\Sigma$ W4A6 precision and Hunyuan-DiT W4A8 precision.
Model Method Prec. Preference(%) $\uparrow$
@@
FP4DiT W4A6 38.81
Table 4: AdaRound/BRECQ user preference study between FP4DiT, Q-Diffusion and TFMQ-DM on Hunyuan DiT with W4A8 and W4A6 quantization.
Our first test focused on comparing FP4DiT to other methods explicitly designed for DiT PTQ: ViDiT-Q and Q-DiT. This study consists of 120 prompts and 17 human participants. We compare generated image quality for PixArt-$\Sigma$ at W4A6 precision and Hunyuan at W4A8 precision. Table 4 reports our findings. FP4DiT is preferred over 50% of the time for PixArt-$\Sigma$ W4A6 precision level, followed by Q-DiT and then ViDiT-Q. This result aligns with Table 2 where FP4DiT achieves the best result and Q-DiT outperforms ViDiT-Q in this setting. For Hunyuan at W4A8 the preference vote share for baseline methods grow, but they do not overtake FP4DiT, which demonstrates the veracity of our approach.
Next, we conduct an additional study comparing FP4DiT to baseline approaches originally designed with U-Net DMs in mind but which utilize advanced PTQ via AdaRound and BRECQ: Q-Diffusion and TFMQ-DM. Specifically, we consider 75 random prompts and 15 human participants. The baseline DM is Hunyuan quantized to either W4A8 or W4A6 precision. Table 4 reports our findings. At A8 precision FP4DiT clearly outperforms the other methods with a majority of the images being favored, while it also obtains almost 40% of the votes at A6 precision.
4.3 Ablation Studies
We conduct ablation studies on the PixArt-$\alpha$ model to verify the contribution of each component of FP4DiT. The experiment settings are consistent with Section 4.2 unless specified. Additional ablation studies can be found in the appendix.
4.3.1 Effect of FP Format.
@@ 4.3.3 Effect of Scale-Aware AdaRound
Figure 10: The performance of different AdaRound methods with different calibration budgets on W4A16 PixArt-$\alpha$ quantization generating 1k images. The stars indicate the optimal budget for INT and scale-aware AdaRound.
To further verify our scale-aware AdaRound for FP quantization, we compare our scale-aware AdaRound to the origin AdaRound with INT quantization in the W4A16 setting. Note that the calibration budget is crucial for the performance of weight reconstruction. BRECQ (Li et al., 2021) uses 20k as default and this setting is inherited by prior DM quantization research like Q-Diffusion and TFMQ-DM. Thus, we configured calibration budgets at {2.5k, 5k, 10k, 20k} to ensure a fair comparison. Figure 10 outlines the CLIP of different AdaRound methods on PixArt-$\alpha$ under different calibration budgets. Scale-aware AdaRound achieving optimal budget with 8 times fewer calibration steps than INT AdaRound, highlights its effectiveness in reducing calibration costs without compromising reconstruction quality.
@@ 4.4 Hardware Cost Comparison
Finally, we measure the hardware latency impact of FPQ compared to more traditional INT quantization and the W16A16. Specifically, we consider the quantization latency cost to execute some of the common, yet costlier weight/activation operations in a DiT, namely the self-attention mechanism and feedforward module. Specifically, the linear layers corresponding to the self-attention mechanism (Q, K, V and output projection) all share the same weight/activation dimensions, while the feedforward consists of two linear layers which expand and then contract the token/patch embedding dimension, respectively. Additionally, weights and activations may not share the same bit precision, imposing additional dequantization cost.
We develop a CUDA 12.8 kernel and measure hardware latency on an Nvidia 5080, a Blackwell GPU which supports low-bit FP formats. We consider the self-attention weight/activation dimensions found in PixArt-$\alpha$/$\Sigma$ when generating an $1024\times 1024$ image. Table 6 reports our findings. The result indicates that, despite the general expectation that floating-point computations are more demanding than integer operations, the latency results are nearly identical across all tested quantization precisions. Specifically, although A6 precision imposes some additional overhead compared to A8 (either for FP or INT), this is quite minor, leading to an overall speedup around 1.5x compared to W16A16, much like the popular W4A8 integer quantization (Li et al., 2023; Zhao et al., 2024; Chen et al., 2025).
Leveraging this finding, our FPQ method can achieve superior performance without sacrificing computational efficiency. In fact, this latency results align with the stated computation rate of different INT and FP formats across different Nvidia GPUs, which we report in Table 7. Specifically, GPUs can handle INT and FP quantization with similar computational throughput, which ensures that FP-based quantization does not introduce additional computational overhead. Lastly, although FP6 shares the same compute throughput as 8-bit formats (see Table 7), it brings memory saving (25% smaller tensor) which is essential for low-bit quantization, as transformers often become memory-bound (Xia et al., 2024) during token generation. As a result, FP6 enhances DiT’s inference efficiency.
We also compare the hardware cost of the quantized FP4DiT model with the full-precision model in terms of memory and energy consumption using two metrics: model size and Bit-Ops (BOPs) (He et al., 2024). Specifically, model size refers to the disk space required to store the model checkpoint weights and scales, while BOPs is a quantization-aware extension of the MACs (Chu et al., 2022; Mills et al., 2023) metric which measures the compute cost of a neural network forward pass. As shown in Table 8, group-wise weight quantizationintroduces only moderate overhead; nonetheless, our method still substantially reduces the model size and BOPs.
Layer W4A8-FP (ms) W4A6-FP (ms) W4A8-INT (ms) W16A16 (ms)
Feedforward Layer 1 303.42 (1.54×) 313.54 (1.49×) 302.56 (1.54×) 465.99
Feedforward Layer 2 313.31 (1.50×) 318.67 (1.47×) 312.80 (1.50×) 469.91
Self-Attention QKV Proj. 78.33 (1.51×) 79.94 (1.48×) 78.16 (1.51×) 118.28
Table 6: The latency and the speedup of different linear layers in PixArt-$\alpha$/$\Sigma$ under different quantization precision generating a $1024\times 1024$ image. Result measured on RTX 5080 GPU using CUDA 12.8.
GPU INT8 FP8 FP6 FP4
RTX 4090 (NVIDIA, 2024a) 660.6 TOPS 660.6 TFLOPS
H100 (NVIDIA, 2024c) 1979 TOPS 1979 TFLOPS
RTX 5090 (NVIDIA, 2024a) 838 TOPS 838 TFLOPS 838 TFLOPS 1676 TFLOPS
HGX B100 (NVIDIA, 2024b) 56 POPS 56 PFLOPS 56 PFLOPS 112 PFLOPS
HGX B200 (NVIDIA, 2024b) 72 POPS 72 PFLOPS 72 PFLOPS 144 PFLOPS
Table 7: GPU throughput rates for different low-bit datatype formats. Horizontal line demarcates older Ada Lovelace/Hopper GPUs from state-of-the-art Blackwell series. Older series do not support FP6 and FP4.
@@
Table 8: The comparison of model size and Bit-Ops of different Precision on PixArt-$\alpha$. G128 denotes group-wise weight quantization with a group size of 128. Lower is better.